OpenStudy (anonymous):

PLEASE PLEASE Help ME: Trevor is tiling his bathroom floor, which has an area that is represented as 175r^5 square inches. Each tile has an area of sqrt64r^17. The total number of tiles used can be represented by the expression below. 175r^5/sqrt64r^17 Simplify the expression for the TOTAL NUMBER of tiles used. Show your work PLEASE. You're AWESOME.

OpenStudy (anonymous):

Does sqrt64r^17 imply \(\sqrt{64} r^{17}\) or \(\sqrt{64 r^{17}}\)?

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Hint:\[\Large \sqrt{64 r^{17}}=\sqrt{64}\sqrt{r^{17}}=8\left( r^{17} \right)^{\frac{1}{2}}=8r^{\frac{17}{2}}\]